### v2

@@ Abstract
Krylov Method, Randomized algorithms, Matrix Functions, Partial Differential Equations
@@ 1 Introduction
Most applications are only interested in the action of $f(\mathbf{A})$ over a vector $\mathbf{b}\in\mathbb{C}^{n}$. Since explicitly forming the full matrix $f(\mathbf{A})$ is unfeasible for a large $\mathbf{A}$, they generally employ Krylov subspace methods [22, 23] to approximate $f(\mathbf{A})\mathbf{b}$ directly. The key ingredient for these methods is the Arnoldi process that constructs an orthonormal basis for the Krylov subspace $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$. However, the Krylov approximation for $f(\mathbf{A})\mathbf{b}$ requires that the entire basis be stored in memory, limiting the size of the problem that can be solved. Furthermore, for non-Hermitian matrices, the cost of orthogonalization can quickly become overwhelming as it grows quadratically with the basis size $m$.
@@
In this paper, we propose a new randomized algorithm for accelerating restarted Krylov methods. In essence, we replace the standard Arnoldi procedure with the randomized version [39] to quickly construct a non-orthogonal, but well-conditioned Krylov basis at each restart cycle. We show that this modification significantly improves the performance of the restarted methods while maintaining the same level of accuracy and stability. Randomization has already been considered before in [13, 24] as a way to accelerate Krylov methods for evaluating general matrix functions. However, none explore restarting procedures.
Another contribution of this paper is to analyse the performance and stability of randomized methods for computing matrix functions that arise from finite element problems that are commonly found in many scientific applications. Although randomized methods behave rather well in moderately conditioned and small problems [13, 24], it is still unknown how they will act when facing a very large, sparse and ill-conditioned problem, which is the typical case when dealing with discretization matrices that arise from finite element methods.
The rest of the paper is organized as follows. Section 2 reviews the theoretical background for standard and randomized Krylov approximation. Section 3 presents other approaches for accelerating Krylov methods using randomization. Section 4 describes the restarting procedure for the randomized Krylov method. Section 5 illustrates the performance and convergence of the proposed method and compares it against the state-of-the-art. In Section 6, we conclude our paper.
@@ 2.1 Notation
2.2 The FOM approximation for matrix functions
@@
The columns of $\mathbf{V}_{m}$ spans $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$ and are ordered such that $\mathbf{v}_{1}=\mathbf{b}/\beta$ with $\beta=\|\mathbf{b}\|$. While $\mathbf{H}_{m}=\mathbf{V}_{m}^{*}\mathbf{A}\mathbf{V}_{m}\in\mathbb{C}^{m\times m}$ is an upper Hessenberg matrix representing the compression of $\mathbf{A}$ onto $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$ with respect to the basis $\mathbf{W}_{m}$. The FOM approximation for $f(\mathbf{A})\mathbf{b}$ is then defined as
@@ 4 Restarted Randomized Krylov
The convergence of restarted methods has only been established for entire functions of order $1$ [14, Theorem 4.2] and Stieljes functions [18], while the general case remains an open problem. With randomization, a rigorous analysis become even more challenging. For this reason, we analyse the convergence of the randomized methods based solely on numerical examples.

### v3

@@ Abstract
Krylov Method, Randomized algorithms, Matrix Functions, Partial Differential Equations, Finite Element Method
@@ 1 Introduction
Most applications are only interested in the action of $f(\mathbf{A})$ over a vector $\mathbf{b}\in\mathbb{C}^{n}$. Since explicitly forming the full matrix $f(\mathbf{A})$ is unfeasible for a large matrices, they generally employ Krylov subspace methods [22, 23] to approximate $f(\mathbf{A})\mathbf{b}$ directly. The key ingredient for these methods is the Arnoldi process that constructs an orthonormal basis for the Krylov subspace $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$. However, the evaluation of $f(\mathbf{A})\mathbf{b}$ requires that the entire basis be stored in memory, limiting the size of the problem that can be solved. Furthermore, for non-Hermitian matrices, the cost of orthogonalization can quickly become overwhelming as it grows quadratically with the basis size $m$.
@@
In this paper, we propose a new randomized algorithm for accelerating restarted Krylov methods. In essence, we replace the standard Arnoldi procedure with the randomized version [39] to quickly construct a non-orthogonal, but well-conditioned Krylov basis at each restart cycle. We show that this modification significantly improves the performance of the restarted methods while maintaining the same level of accuracy and stability. In fact, the convergence rate of the randomized method is often superior than the classical version in most of our tests. Randomization has already been considered before in [13, 24] as a way to accelerate Krylov methods for evaluating general matrix functions. However, restarting procedures remains unexplored.
Although randomized methods behave rather well in moderately conditioned and small problems [13, 24], it is still unknown how they will act when facing badly conditioned matrices that are commonly found in many real-world applications. In particular, the discretization of PDEs via finite element method (FEM) is well-known to produce very large, sparse and ill-conditioned matrices. Therefore, our second contribution is an extensive study of the convergence and performance of several randomized methods through a set of large-scale numerical experiments, focusing exclusively on the finite element problems.
The rest of the paper is organized as follows. Section 2 reviews the background theory of the Krylov subspace methods. It also describes the randomized Arnoldi iteration and how it can be integrated into Krylov methods. Section 3 presents other approaches for accelerating Krylov methods using randomization. Section 4 describes the restarting procedure for the randomized Krylov method. Section 5 illustrates the performance and convergence of the proposed method and compares it against the state-of-the-art. In Section 6, we conclude our paper.
@@ 2.1 Notation
2.2 Krylov Subspace Methods
@@
The columns of $\mathbf{V}_{m}$ spans $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$ and are ordered such that $\mathbf{v}_{1}=\mathbf{b}/\beta$ with $\beta=\|\mathbf{b}\|$. While $\mathbf{H}_{m}=\mathbf{V}_{m}^{*}\mathbf{A}\mathbf{V}_{m}\in\mathbb{C}^{m\times m}$ is an upper Hessenberg matrix representing the compression of $\mathbf{A}$ onto $\mathcal{K}_{m}(\mathbf{A},\mathbf{b})$ with respect to the basis $\mathbf{V}_{m}$. The FOM approximation for $f(\mathbf{A})\mathbf{b}$ is then defined as
@@ 4 Restarted Randomized Krylov
The convergence of restarted methods has only been established for entire functions of order $1$ [14, Theorem 4.2] and Stieljes functions [18], while the general case remains an open problem. With randomization, a rigorous analysis becomes even more challenging. For this reason, we analyse the convergence of the randomized methods based solely on numerical examples.